One month Lisa rented
2
movies and
3
video games for a total of
$25
. The next month she rented
4
movies and
8
video games for a total of
$61
. Find the rental cost for each movie and each video game.

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x = price movie
y = price game
---
2x + 3y = 25
4x + 8y = 61
---
put the system of linear equations into standard form
---
2x + 3y = 25
4x + 8y = 61
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = price movie = 4.25
y = price game = 5.50
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

One month Lisa rented 2 movies and 3 video games for a total of $25
The next month she rented 4 movies and 8 video games for a total of $61
. Find the rental cost for each movie and each video game.
----------------
Equations:
2m + 3v = 25
4m + 8v = 61
-----
Modify for elimination::
8m + 12v = 100
8m + 16v = 122
-----
Subtract and solve for "v"::
4v = 22
v = $5.50 (cost of a video game)
---
Solve for "m":
2m + 3v = 25
2m + 3*5.5 = 25
2m + 16.5 = 25
2m = 8.50
m = $4.25 (cost of a movie)
-------------
Cheers,
Stan H.
------------------